A ball is tied at the end of a rope and swings in a circle at a constant speed, doing 60 revolutions a minute. What is the frequency, which period does this movement have?
1
Expert's answer
2020-09-14T10:20:43-0400

The frequency of rotation, "f" means how many rotations take place in a certain amount of time.


"f=\frac{60}{1} min^{-1} =\frac{60}{60} s^{-1}=1s^{-1}"


Period, T:

"T=\frac{1}{f} \\ 
T=1 s \\
Answer: f=1s^{-1}, T=1s"